The radius of the outer spear of a spherical capacitor is five times the radius of its inner shell . What are the dimensions of these capacitor if its capacitance is 5.00 picofarad?


1
Expert's answer
2022-03-16T09:41:50-0400

The capacitance is given by


"C=4\pi\epsilon_0(R_2*R_1)/(R_2-R_1)""R_2=5R_1"

So


"5.00*10^{-12}=4*3.14*8.85*10^{-12}\frac{5R_1^2}{4R_1}""R_1=0.0360\:\rm m=3.60\:\rm cm""R_2=5R_1=18.0\:\rm cm"